A. Balance the following equations by adding coefficients. Do not leave blank spaces - use a 1 if necessary. Identify the tB. Complete the following equations by writing the formula(s) of the product(s) and balancing the equation with coefficients.C. Write the balanced chemical equation for the following chemical reactions. Include the correct formulas for all reactants20. Copper (II) sulfate reacts with aluminum to form aluminum sulfate and copper. D. Complete the following stoichiometry proc. How many grams of Cr metal would be needed to produce 38.2g of chromium (III) oxide? Sulfuric acid (H2SO4) is a component23. Pentane combusts with oxygen to form carbon dioxide and water by the following reaction: C5H12(1) + 8 O2(g) → 5 CO2(g) +624. Hydrogen reacts with nitrogen to form ammonia by the following equation: 3 H2(g) + N2(g) → 2 NH3(g) a. What is the limitic. Calculate the percent yield of ammonia when 7.45 grams of ammonia was experimentally obtained from the reaction? d. How mu
Answer #1

A. 1. HYAS O7 As2O5HoO

There are 7 O on left and 6 O on right. So to balance it out write two as coefficient of H2O.

H_4As_2O_7\rightarrow As_2O_5+2H_2O

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of H4As2O7 and As2O5.

1H_4As_2O_7\rightarrow 1As_2O_5+2H_2O

This is the balanced chemical equation. As one reactant (H4As2O7) breaks down to form two products (As2O5 and H2O) so it is a decomposition reaction.

2. N_2+O_2 \rightarrow N_2O

There are 2 O on left and 1 O on right. So to balance it out write two as coefficient of N2O.

N_2+O_2 \rightarrow 2N_2O

There are 2 N on left and 4 N on right, so to balance it out write 2 as coefficient of N2

2N_2+O_2 \rightarrow 2N_2O

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of O2.

2N_2+1O_2 \rightarrow 2N_2O

This is the balanced chemical equation. As two reactants (N2 and O2) combine to form one product (N2O) so it is a combination reaction.

3. NaI+Br_2 \rightarrow NaBr+I_2

There are 2 Br on left and 1 Br on right. Also there are 2 I on right and one on left. So to balance these out write two as coefficient of NaI and NaBr.

2NaI+Br_2 \rightarrow 2NaBr+I_2

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of I2 and Br2.

2NaI+1Br_2 \rightarrow 2NaBr+1I_2

This is the balanced chemical equation. As one reactive element (Br2 in this case) replaces the less reactive element (I2 in this case) from its compound (NaI in this case), so this is a single replacement reaction.

4. PbCrO_4+HNO_3 \rightarrow Pb(NO_3)_2+H_2CrO_4

There are 2 H on right and 1 H on left. There are 2 NO3 groups on right and one on left. So to balance these out write two as coefficient of HNO3.

PbCrO_4+2HNO_3 \rightarrow Pb(NO_3)_2+H_2CrO_4

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of PbCrO4 and H2CrO4.

1PbCrO_4+2HNO_3 \rightarrow Pb(NO_3)_2+1H_2CrO_4

This is the balanced chemical equation. As there is a mutual exchange of ions between the two reacting compounds to from two new compounds this is a double replacement reaction.

5. C_3H_8+O_2 \rightarrow CO_2+H_2O

There are 3 C on left and one on right. So to balance it out write three as coefficient of CO2.

C_3H_8+O_2 \rightarrow 3CO_2+H_2O

There are 8 H on left and 2 on right so to balance it out write 4 as coefficient of H2O.

C_3H_8+O_2 \rightarrow 3CO_2+4H_2O

Now there are 3x2+4=6+4=10 O on right and 2 O on left. So to balance it out write 5 as coefficient of O2

C_3H_8+5O_2 \rightarrow 3CO_2+4H_2O

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of C3H8

1C_3H_8+5O_2 \rightarrow 3CO_2+4H_2O

This is the balanced chemical equation. As in this reaction a hydrocarbon (C3H8) reacts with oxygen (O2) to form CO2 and H2O, this is a combustion reaction.

6. TiCl_4+Mg \rightarrow MgCl_2+Ti

There are 4 Cl on left and 2 Cl on right. So to balance it out write two as coefficient of MgCl2.

TiCl_4+Mg \rightarrow 2MgCl_2+Ti

There are 2 Mg on right and one on left, so to balance it out write 2 as coefficient of Mg.

TiCl_4+2Mg \rightarrow 2MgCl_2+Ti

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of TiCl4 and Ti.

1TiCl_4+2Mg \rightarrow 2MgCl_2+1Ti

This is the balanced chemical equation. As one reactive element (Mg in this case) replaces the less reactive element (Ti in this case) from its compound (TiCl4 in this case), so this is a single replacement reaction.

7. CuSO_4+KCN \rightarrow Cu(CN)_2+K_2SO_4

There are 2 CN on right and 1 on left. So to balance it out write two as coefficient of KCN.

CuSO_4+2KCN \rightarrow Cu(CN)_2+K_2SO_4

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of CuSO4 and K2SO4.

1CuSO_4+2KCN \rightarrow Cu(CN)_2+1K_2SO_4

This is the balanced chemical equation. As there is a mutual exchange of ions between the two reacting compounds to from two new compounds this is a double replacement reaction.

8. Ca(ClO_3)_2 \rightarrow CaCl_2+O_2

There are 3x2=6 O on left and 2 O on right. So to balance it out write three as coefficient of O2.

Ca(ClO_3)_2 \rightarrow CaCl_2+3O_2

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of Ca(ClO3)2 and CaCl2.

1Ca(ClO_3)_2 \rightarrow 1CaCl_2+3O_2

This is the balanced chemical equation. As one reactant (Ca(ClO3)2) breaks down to form two products (CaCl2 and O2) so it is a decomposition reaction.

9. Na_2O +H_2O \rightarrow NaOH

There are 2 Na on left and 1 O on right. So to balance it out write two as coefficient of NaOH.

Na_2O +H_2O \rightarrow 2NaOH

Now we can see that all the atoms are balanced in this reaction. Write 1 as coefficient of Na2O and H2O.

1Na_2O +1H_2O \rightarrow 2NaOH

This is the balanced chemical equation. As two reactants (Na2O and H2O) combine to form one product (NaOH) so it is a combination reaction.

10. C_6H_{14}+O_2 \rightarrow CO_2+H_2O

There are 6 C on left and one on right. So to balance it out write six as coefficient of CO2.

C_6H_{14}+O_2 \rightarrow 6CO_2+H_2O

There are 14 H on left and 2 on right so to balance it out write 7 as coefficient of H2O.

C_6H_{14}+O_2 \rightarrow 6CO_2+7H_2O

Now there are 6x2+7=12+7=19 O on right and 2 O on left. So to balance it out write 19/2 as coefficient of O2

C_6H_{14}+\frac{19}{2}O_2 \rightarrow 6CO_2+7H_2O

Multiply the whole equation by two

2C_6H_{14}+19O_2 \rightarrow 12CO_2+14H_2O

Now we can see that all the atoms are balanced in this reaction.

This is the balanced chemical equation. As in this reaction a hydrocarbon (C6H14) reacts with oxygen (O2) to form CO2 and H2O, this is a combustion reaction.